# mybibliography.bib

### v1

@@ 5 Discussion
We introduce a regularization scheme for recurrent neural nets influenced by the underlying stable dynamics of the recurrent system. We do this by embedding the recurrent neurons into a fixed metric space, and differentially penalizing weights according to their distance on this metric space. By embedding neurons into manifolds, we reparameterized their dynamics into sparse, geometrically organized structures. We then implemented our regularizer on navigation and natural language processing recurrent architectures, showing that it functions as a highly effective method of sparsification.
The resulting sparse networks have several features unusual in sparse neural nets. Randomly sparsified RNNs have a weakly Ramanujan sparse structure [friedman2003relative], meaning they are highly connected despite being sparse. This is a desirable property in network theory [hoory2006expander], but not a feature of networks sparsified with moduli regularization. Networks trained with moduli regularizers have highly correlated adjacency structures, reflecting that their connections are drawn from continuous metric spaces.
Our method is distinguished from previous sparsification methods in that it creates stable sparse models in the sense of the lottery ticket hypothesis [frankle2018lottery, frankle2020linear]. Prior work in model sparsification has found that training sparse models is heavily dependent on the weight initialization [frankle2018lottery, gale2019state]. Incorporating moduli regularization into sparsification alleviates these problems in some examples. In addition to creating superior sparse models, the underlying sparse architectures are stable: new models, with randomized initializations, can be trained on previously learned sparse architectures to high fidelity. This opens the door to future work analyzing the limitations of different sparse architectures, and the possibility of training a priori sparse models. Additionally, moduli regularization can be used as an exploratory technique to identify plausible candidates for manifolds underlying neural computations.
In future work, we are particularly interested in studying multi-modular regularization systems, in which different neurons are embedded in different moduli spaces, as well as adapting our work to different neural architectures. An important generalization of our work is a method which better combines manifold learning with moduli regularization to dynamically choose optimal moduli spaces for a given network.
@@ A.1 Navigation RNN
We use the same hyperparameters and architecture as [sorscher2023unified], as listed below in Table 6, with the exception of increasing the batch size and decreasing the number of training epochs. We also remove the $L_{2}$ regularization that was applied in [sorscher2023unified], since it substantially slows training. The code is freely available.
@@
Optimizer Adam [kingma2014adam]
@@
Loss Cross entropy [mannor2005cross] of softmax of Gaussians
@@
The input and output functions of the RNN and the focus on landmarks is not as strange as it might appear. A very similar, though much larger, architecture was used in [mirowski2018streetnav] for navigation in large cities based purely on visual input. The rationale is that introduction of new landmarks is arbitrarily extensible, lending it greater potential generalizability of the system.
@@ A.2 NLP RNN
We use the recommended hyperparameters of and architecture of [pytorch_examples], as listed below in Table 7. The code is available under the BSD 3-Clause License. For training and testing data, we use the Wikitext-2 dataset of high quality Wikipedia articles [merity2016pointer], available under Creative Commons Attribution Share Alike 3.0 license.
@@
The data was split as in the Wikitext-2 dataset [merity2016pointer], which corresponds to roughly 82% of data reserved for training, training, 8% for validation, and 10% for test. All reported results are on the test dataset.
@@
We additionally ran exploratory experiments to identify good thresholds for sparsification, for which we did not record compute time. The interested reader is recommended Tu’s (2011) book [tu2011manifolds].
@@ Appendix B Manifolds
A manifold is a topological space which locally is isomorphic to the Euclidean space [hatcher2005algebraic]. A sphere like the earth’s surface, for example, has global curvature: it isn’t isomorphic to $\mathbb{R}^{2}$. However, a person walking on the earth could easily mistake it as flat: a square grid is a fine way to organize something as large as a city. This motivates the following definitions:
@@
We particularly care about compact manifolds. For manifolds that are embedded in Euclidean space (for example, a sphere or torus in three dimensional space), these are spaces which contain all their limit points and are bounded. Compact spaces have the nice property that there is an upper bound on how far away pairs of points can be.
@@ Appendix C Covering maps
is a covering space.
@@
As a result, there is a covering map $p\times p:\mathbb{R}^{2}\to S^{1}\times S^{1}$ from two dimensional Euclidean space to the torus. There is a special kind of covering map, called a universal covering map. This is a covering space $X\to Y$ where $X$ has trivial fundamental group: that is, all loops in the space can be filled in. The interested reader is recommended Hatcher’s (2005) book [hatcher2005algebraic] for more details.
@@ Appendix E Inhibitor functions
Results are shown in Tables 8 and 9 for navigation and natural language processing, respectively.
@@
Table 9: Accuracy (standard deviation) of next-word predictor RNNs (Section 4.2). The difference of Gaussians inhibitor provides substantially better results, both during sparsification and in lottery ticket experiments.
@@
Baseline 8.02% (1.73) 8.72% (1.11) 2.36% (0.15)
90% Sparse 10.31% (1.33) 7.41% (3.78) 4.22% (2.87)
90% Lottery 9.77% (0.81) 6.62% (3.32) 2.85% (0.00)
98% Sparse 9.59% (1.07) 10.08% (1.01) 4.34% (3.29)
98% Lottery 9.17% (0.51) 7.62% (0.96) 2.81% (0.07)
@@ Appendix G Sensitivity to the weight decay factor
The influence of regularization on trained models is controlled by the weight decay constant it is multiplied by ($\lambda$ in Equation (4)). We show the results of varying the regularizing coefficient $\lambda$ for our navigation experiments in Table 10 and the regularizing coefficient $\lambda$ for NLP in Table 11.
Table 10: We show sensitivity of the sparsification of our navigation experiments at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report mean error (cm) of a single trial.
$1$ $10^{-1}$ $10^{-2}$ $10^{-3}$
$L_{1}$ 90% Sparse 92.26 92.26 86.53 90.83
$L_{1}$ Lottery 54.55 54.55 45.78 68.31
Torus 90% Sparse 5.22 5.70 5.54 5.75
Torus Lottery 79.70 13.03 30.51 51.80
Shuffled 90% Sparse 58.38 20.66 5.51 85.78
Shuffled Lottery 86.13 32.93 42.69 58.62
Table 11: We show sensitivity of the sparsification of our NLP experiment at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report accuracy of the next-word prediction of a single trial.
1 $10^{-1}$ $10^{-2}$ $10^{-3}$
$L_{1}$ 90% Sparse 2.26% 2.85% 2.85% 1.20%
$L_{1}$ Lottery 2.26% 2.85% 2.85% 2.85%
Torus 90% Sparse 9.18% 9.94% 2.85% 2.85%
Torus Lottery 10.26% 9.82% 2.44% 2.59%
Shuffled 90% Sparse 8.32% 8.84% 8.99% 2.26%
Shuffled Lottery 9.52% 8.72% 2.85% 2.85%
@@ Appendix H Trained embeddings to learn manifolds
To test for manifold learning with trained embeddings, we apply persistence homology [edelsbrunner2008persistent] to navigation experiments (Section 4.1) with moduli space $\mathbb{R}^{3}$ and difference of Gaussians inhibitor function (Equation (7)) shown in Table 3. The persistence diagrams for three of the neural point cloud are shown in Figure 4. There is no clear learning of a manifold present: the neurons appear randomly distributed. For these experiments, we initialized neurons on a uniform distribution in $[0,10]^{3}\subset\mathbb{R}^{3}$.

### v2

@@ 5 Discussion
We introduce a regularization scheme for recurrent neural nets influenced by the underlying stable dynamics of the recurrent system, which can be initialized by the researcher by applying special domain-specific knowledge. We do this by embedding the recurrent neurons into a fixed metric space, and differentially penalizing weights according to their distance on this metric space. By embedding neurons into manifolds, we reparameterized their dynamics into sparse, geometrically organized structures. We then implemented our regularizer on navigation, natural language processing, and adding problem recurrent architectures, showing that it functions as a highly effective method of sparsification.
Our method is distinguished from previous sparsification methods in that it creates relatively stable sparse models in the sense of the lottery ticket hypothesis [9, 10]. Prior work in model sparsification has found that training sparse models is heavily dependent on the weight initialization [9, 11]. Incorporating moduli regularization into sparsification alleviates these problems in some examples. In addition to creating superior sparse models, the underlying sparse architectures are stable: new models, with randomized initializations, can be trained on previously learned sparse architectures to high fidelity. This opens the door to future work analyzing the limitations of different sparse architectures, and the possibility of training a priori sparse models. Additionally, moduli regularization can be used as an exploratory technique to identify plausible candidates for manifolds underlying neural computations.
A particularly interesting result of our experiments is that we demonstrate the importance of global weight organization in RNNs. The results of Table 3 are particularly incisive: the RNN is able to be effectively sparsified exactly when the moduli space is a good approximation of the underlying dynamics.
In future work, we are particularly interested in studying multi-modular regularization systems, in which different neurons are embedded in different moduli spaces, as well as adapting our work to different neural architectures and use-cases. An important generalization of our work is a method which better combines manifold learning (such as [7, 17, 24, 42, 48]) with moduli regularization to dynamically choose optimal moduli spaces for a given network.
@@ A.1 Navigation RNN
We use the same hyperparameters and architecture as [49], as listed below in Table 11, with the exception of increasing the batch size and decreasing the number of training epochs. We also remove the $L_{2}$ regularization that was applied in [49], since it substantially slows training. The code is freely available.
@@
Optimizer Adam [25]
@@
Loss Cross entropy [32] of softmax of Gaussians
@@
$L_{1}$ regularizers used the same regularizing factor as moduli regularizers.
@@
The input and output functions of the RNN and the focus on landmarks is not as strange as it might appear. A very similar, though much larger, architecture was used in [34] for navigation in large cities based purely on visual input. The rationale is that introduction of new landmarks is arbitrarily extensible, lending it greater potential generalizability of the system. We depict a diagram of the full architecture in Figure 4, and in shorter form in Figure 2.
@@ A.2 NLP RNN
We use the recommended hyperparameters of and architecture of [3], as listed below in Table 12. The code is available under the BSD 3-Clause License. For training and testing data, we use the Wikitext-2 dataset of high quality Wikipedia articles [33], available under Creative Commons Attribution Share Alike 3.0 license.
@@
$L_{1}$ regularizers used the same regularizing factor as moduli regularizers.
The data was split as in the Wikitext-2 dataset [33], which corresponds to roughly 82% of data reserved for training, training, 8% for validation, and 10% for test. All reported results are on the test dataset.
@@
We additionally ran exploratory experiments to identify good thresholds for sparsification, for which we did not record compute time. The interested reader is recommended Tu’s (2011) book [52].
A.3 Adding problem RNN
We used the hyperparameters shown in Table 13.
Table 13: Hyperparameters for the adding problem experiments (Section 4.3).
Hyperparameters Values
Optimizer Adam [25]
Batch size 32
Batches tested 500,000
Sequence length 50, or as specified
Learning rate $10^{-4}$
Hidden dim 128
Bias Yes
Nonlinearity ReLU
Regularizing factor $10^{-3}$
Loss Mean square error
Grad clip 0.5
$L_{1}$ regularizers used the same regularizing factor as moduli regularizers.
A.4 Speech recognition
We use the same hyperparameters as [39], as shown in Table 14, with the exception of our use of a single layer RNN, rather than a multi-layer bidirectional GRU.
Table 14: Hyperparameters for speech recognition experiments.
Hyperparameters Values
Optimizer AdamW [29]
Nonlinearity tanh (recurrent), GELU [21] (non-recurrent)
Number of hidden layers 1
Bias Yes
Number of convolutional layers 3
Convolutional channels 32
Convolutional kernel $3\times 3$, with stride 1
Batch size 20
Hidden dim 512
Dropout 0.1
Learning rate $5\times 10^{-4}$
Learning rate annealing strategy Linear
Epochs 10
Regularizing constant $10^{-2}$
Loss CTCLoss [16]
$L_{1}$ regularizers used the same regularizing factor as moduli regularizers.
The data was split as follows: 80% training, 20% testing. The Librispeech dataset [41] is available under the Creative Commons 4.0 license.
Based on the Deep Speech 2 model [39], the RNN structure is given by:
  1. 1.
    Compute the Mel spectogram [46] of the data
  2. 2.
    Three residual convolutional layers process the spectogram
  3. 3.
    Three bidirectional RNNs process the output of the convolutional layers
  4. 4.
    Two linear layers with GELU nonlinearity
We then used the Connectionist Temporal Classification Loss (CTCLoss) [16], a paradigm designed for variable many-to-one mappings, as a loss function.
@@ Appendix B Manifolds
A manifold is a topological space which locally is isomorphic to the Euclidean space [20]. A sphere like the earth’s surface, for example, has global curvature: it isn’t isomorphic to $\mathbb{R}^{2}$. However, a person walking on the earth could easily mistake it as flat: a square grid is a fine way to organize something as large as a city. This motivates the following definitions:
@@
A manifold is $\mathcal{C}^{k}$ if, for each pair of elements of the open cover $U_{i},U_{j}$, with $f_{i}:U_{i}\to\mathbb{R}^{n}$, $f_{j}:U_{j}\to\mathbb{R}^{n}$ homemorphisms, the functions $f_{i}\circ f_{j}^{-1}$ and $f_{j}\circ f_{i}^{-1}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ are $k$-times differentiable.
We particularly care about compact manifolds. For manifolds that are embedded in Euclidean space (for example, a sphere or torus in three dimensional space), these are spaces which contain all their limit points and are bounded. Compact spaces have the nice property that there is an upper bound on how far away pairs of points can be. Common examples of manifolds include the circle, the sphere, and the torus. A figure 8, however, is not a manifold: at the intersection point of the two loops it is not homeomorphic to the line.
@@ Appendix C Covering maps
is a covering space. We illustrate this example in Figure 5.
Figure 5: $\mathbb{R}^{1}$, represented as a blue line, is a covering space for $S^{1}$, represented as the black circle. The projection map is $x\mapsto(\cos x,\sin x)$.
@@
As a result, there is a covering map $p\times p:\mathbb{R}^{2}\to S^{1}\times S^{1}$ from two dimensional Euclidean space to the torus. There is a special kind of covering map, called a universal covering map. This is a covering space $X\to Y$ where $X$ has trivial fundamental group: that is, all loops in the space can be filled in. The interested reader is recommended Hatcher’s (2005) book [20] for more details.
@@ Appendix E Inhibitor functions
Results are shown in Tables 15 and 16 for navigation and natural language processing, respectively. We illustrate how these inhibitor functions appear in Figure 6.
Figure 6: We illustrate the manner in which different moduli spaces and different inhibitor functions induce different weights in the hidden update matrix. A neuron $y$ is fixed. Red points depict output neurons $z$ with low regularizing values, and therefore potentially large weights $w_{yz}$, while blue points are output neurons $z^{\prime}$ with high regularizing values, and therefore smaller potential weights $w_{yz^{\prime}}$. Row (1) depicts the circle, sphere, and Klein bottle, while row (2) depicts the torus. Column (a) has inhibitor function $f(d)=d^{2}$, column (b) has inhibitor function $f(d)=10-10(e^{-\frac{d^{2}}{5}}-e^{-d^{2}})$, and column (c) has inhibitor function $f(d)=10+10\cos(2d)$.
@@
Table 16: Perplexity Accuracy (standard deviation) of next-word predictor RNNs (Section 4.2). The difference of Gaussians inhibitor provides substantially better results, both during sparsification and in lottery ticket experiments.
@@
0% Sparse 69.26 (16.31) 60.49 (11.23) 130.96 (27.59)
90% Sparse 43.41 (7.53) 74.65 (38.75) 112.46 (38.98)
90% Lottery 54.80 (8.70) 81.91 (27.02) 113.77 (0.50)
98% Sparse 55.80 (13.96) 48.75 (10.42) 103.11 (31.08)
98% Lottery 61.44 (6.06) 75.71 (9.19) 114.57 (1.24)
@@ Appendix G Sensitivity to the weight decay factor
The influence of regularization on trained models is controlled by the weight decay constant it is multiplied by ($\lambda$ in Equation (5)). We show the results of varying the regularizing coefficient $\lambda$ for our navigation experiments in Tables 17 and 18 and the regularizing coefficient $\lambda$ for NLP in Tables 19 and 20. Due to the computational costs, we study both cases only at $90\%$ sparsity, and at 10 different weight decay values. Due to a typo during execution, navigation experiments and NLP experiments occurred at slightly different weight decay values—we decided that the results were sufficiently uniform that it was not worth rerunning the experiments for precisely equal comparisons. All results here record different trials than the those presented in the main paper, even when hyperparameters are identical in the two settings.
$1$ $10^{-1}$ $10^{-2}$ $10^{-3}$ $10^{-4}$
$L_{1}$ 90% Sparse 62.37 (12.91) 58.93 (13.95) 60.49 (17.51) 73.93 (9.47) 70.15 (3.17)
Torus 90% Sparse 41.75 (0.49) 9.13 (1.12) 17.88 (6.24) 37.52 (15.22) 58.16 (15.21)
Shuffled 90% Sparse 72.13 (1.72) 23.12 (1.67) 16.92 (5.88) 31.87 (10.94) 70.09 (2.02)
Table 17: We show sensitivity of the sparsification of our navigation experiments at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report mean error (cm) (standard deviation) over five trials.
$6$ $6\times 10^{-1}$ $6\times 10^{-2}$ $6\times 10^{-3}$ $6\times 10^{-4}$
$L_{1}$ 90% Sparse 70.74 (6.52) 60.01 (17.62) 64.73 (18.85) 74.91 (6.08) 57.54 (17.36)
Torus 90% Sparse 47.30 (0.49) 38.16 (1.39) 9.22 (1.92) 12.47 (0.97) 62.66 (17.92)
Shuffled 90% Sparse 58.96 (5.36) 59.58 (7.99) 18.03 (1.99) 15.45 (3.64) 58.69 (14.66)
Table 18: We show extended sensitivity data of the sparsification of our navigation experiments at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report mean error (cm) (standard deviation) over five trials.
1 $10^{-1}$ $10^{-2}$ $10^{-3}$ $10^{-4}$
$L_{1}$ 90% Sparse 106.04 (28.66) 118.17 (3.65) $179.53^{*}$ (99.71) 117.98 (2.73) 101.50 (15.24)
Torus 90% Sparse 114.64 (26.32) 50.46 (12.08) 63.04 (35.18) 90.79 (35.60) 124.77 (16.44)
Shuffled 90% Sparse 85.93 (38.53) 57.14 (18.43) 81.63 (40.31) 133.23 (26.05) 103.54 (27.70)
Table 19: We show sensitivity of the sparsification of our NLP experiment at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report perplexity of the next-word prediction of a single trial. $\,{}^{*}$ marks the exclusion of one trial which suffered representation collapse.
5 $5\times 10^{-1}$ $5\times 10^{-2}$ $5\times 10^{-3}$ $5\times 10^{-4}$
$L_{1}$ 90% Sparse 118.47 (2.62) 106.03 (24.47) 120.29 (2.63) 108.85 (31.30) 184.67 (132.10)
Torus 90% Sparse 111.88 (23.33) 66.47 (21.26) 43.63 (7.80) 103.55 (28.41) 117.96 (1.83)
Shuffled 90% Sparse 110.85 (23.56) 62.17 (13.80) 62.47 (7.05) 162.42 (124.24) 117.39 (16.86)
Table 20: We show extended sensitivity data of the sparsification of our NLP experiment at 90% sparsity of $L_{1}$ regularization, moduli sparsification on the torus, and shuffled moduli regularizers. Here we report perplexity of the next-word prediction of a single trial.
@@ Appendix H Trained embeddings to learn manifolds
To test for manifold learning with trained embeddings, we apply persistence homology [5] to navigation experiments (Section 4.1) with moduli space $\mathbb{R}^{3}$ and difference of Gaussians inhibitor function (Equation (8)) shown in Table 2. The persistence diagrams for three of the neural point cloud are shown in Figure 8. There is no clear learning of a manifold present: the neurons appear randomly distributed. For these experiments, we initialized neurons on a uniform distribution in $[0,10]^{3}\subset\mathbb{R}^{3}$.
@@
Appendix I Moduli regularization for non-recurrent architectures
I.1 Fully connected layers
Given a fully connected layer of a neural net, no changes to Eq. 6 are necessary to create a moduli-like regularizer: the hidden weight update matrix is a special case of a fully connected layer. However, as neurons in layer $k$ of an MLP are not associated with neurons in layer $k+1$, different embeddings from the set of neurons into the manifold $\iota_{k}:[n_{k}]\hookrightarrow\mathcal{M}\hookleftarrow[n_{k+1}]:\iota_{k+1}$ should be chosen.
I.2 Convolutional layers
Regularization and sparsification in convolutional layers of neural nets has historically been unsuccessful [12]. The non-local nature of convolutions causes them to be highly resilient to overfitting, obviating particularly the need for regularization. We describe a generalization of moduli regularization to general convolutional layers in Section I.2.1. Our experiments broadly comported with past results, however: adding moduli regularization to convolutional layers of neural nets results in mildly worse performance, in terms of both training time and validation loss. Nevertheless, this generalization is vital for implementation of moduli regularization in convolutional recurrent neural nets (CRNNS), which we will explore in future work.
The paper [31] implements something very similar: the primary differences to our paper are that they hard-code both weights and sparse structures, rather than regularize their neural nets, and they only consider the Klein bottle and the circle in the second convolutional layer of a neural net. They found some positive results when testing with class correlated noise
I.2.1 Moduli regularization for convolutional layers
For linear/fully connected layers, moduli regularization was implemented by embedding neurons into a manifold $\mathcal{M}$ and penalizing weights between neurons that are very distant. The analogue for convolutional layers is to embed channels into a manifold and penalize weights between channels that are very distant. In equations, the regularization term is
$\sum_{c,c^{\prime}\in\text{Channels}}d_{\mathcal{M}}(i(c),i(c^{\prime}))|T_{c,c^{\prime}}|^{2},$
where $i$ is a chosen embedding of the set of channels into $\mathcal{M}$, $d_{\mathcal{M}}$ is a distance function on $\mathcal{M}$, $T_{c,c^{\prime}}$ is convolved grid between $c$ and $c^{\prime}$ (or 0 if $c^{\prime}$ is not in a layer immediately subsequent to that of $c$), and $|-|^{2}$ is the $L^{2}$ norm on the vector space.